My age is 26, my cousin age is 10. At what age will we be when I am three times his age? What is the equation for that.
Found 3 solutions by Alan3354, ewatrrr, jim_thompson5910:
Answer by Alan3354(69443)

My age is 26, my cousin age is 10. At what age will we be when I am three times his age? What is the equation for that.
-------------
The age difference is 16 years.
3x at 24 & 8
---------------
age + 16 = 3*age

Answer by ewatrrr(24785)

 
Hi,
My age is 26, my cousin age is 10. At what age will we be when I am three times his age?
Let x represent number of years that happens
26 + x = 3(10+x)
-4 = 2x
-2 = x
happened 2 years ago at ages: 24 and 8

Answer by jim_thompson5910(35256)

Let x = number of years

So in x years, you will be 26+x and your cousin will be 10+x

You want to know when you'll be 3 times his age. So ...

Your age in x years = 3*(Cousin's age in x years)

26+x = 3*(10+x)

So the equation is 26+x = 3*(10+x)

-----------------------------------------------------------------------
Let's solve for x


26+x = 3*(10+x)

26+x = 30+3x

26-30 = 3x-x

-4 = 2x

-4/2 = x

-2 = x

x = -2


So because we got a negative number, you will NEVER be 3 times your cousins age after this point in time.

The result of -2 means you WERE 3 times his age 2 years ago since you were 26-2 = 24 yrs old and your cousin was 10-2 = 8 yrs old (notice how 24 = 8*3).
So after this point in time, you'll never be 3 times your cousin's age (since it only happens once).

--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------